Find the equation of the locus of a point so that the square of its distance from (3, -3) is always numerically equal to the slope of the line joining it to the same point.



1
Expert's answer
2021-10-14T12:24:11-0400

"\displaystyle
\text{Consider the point(x,y):}\\
p^2((x,y);(3,-3)) = (x-3)^2 + (y+3)^2\\
\text{The slope of corresponding line equals $\frac{y+3}{x-3}$. As a result, we have an equation:}\\
(x-3)^2 + (y-3)^2=\frac{y+3}{x-3}"